solve.....5y-4+y+18=7y+20-4y
Answer by Jstrasner(112)

Hey,
So for this one, we need to work on each side of the equal sign at a time.
For 5y-4+18
-4+18=14
So the left side equals:
5y+14
Now for the right side of the equals sign:
7y+20-4y
7y-4y=3y
So the right side equals:
3y+20
So the whole equation now equals:
5y+14=3y+20
If you subtract 3y from both sides, the 3y on the right side is cancelled out:
5y-3y+14=3y-3y+20 =>
2y+14=20
Now we subtract 14 from both sides and the 14 on the left side is cancelled out:
2y+14-14=20-14 =>
2y=6
Now divide both sides by 2:
y=3
I hope this helps!